(1+tan1)(1+tan2)................(1+tan45)=2  find k

Grade:12

1 Answers

Gourav
6 years ago
(1+tan1)(1+tan44)=1+tan1+tan44+tan1.tan44[tan (44+1)=tan1+tan44/1-tan1.tan44Or, tan1+tan44=1-tan1.tan44 since tan45=1]Substituting , we get,1+tan1.tan44 +1-tan1.tan44=2Similarly, all the bracket pairs exclusive of (1+tan45) give 2 each. Hence answer=2^23.Therefore k=23